OpenStudy (anonymous):

Square root of -36? Are all negative square roots undefined?

OpenStudy (anonymous):

no because there are imaginary numbers also

OpenStudy (anonymous):



OpenStudy (anonymous):

I haven't learned imaginary numbers!

OpenStudy (jdoe0001):

not just all negative square roots, all EVEN RADICALS with a negative quantity are so-called imaginary \(\bf \large \sqrt[4]{-23}\qquad \sqrt[202]{-79}\qquad \sqrt[24]{-11}\) and so on

OpenStudy (jdoe0001):

check your book, it's surely has such chapter

OpenStudy (anonymous):

I'm doing summer hw, I haven't started the course, so there's no book.

OpenStudy (jdoe0001):

w0t?